Aaron is twice as old as Jack. The sum of their ages three years ago was 45 years. How old are they now?
My problem is when I get the solution, the ages are backwards.
Now Aaron = 2 Now Jack
(Now Aaron - 3) + (Now Jack -3) = 45
Now Aaron + Now Jack = 51
Now Jack = 51 - Now Aaron
Substitute-
Now Aaron = 2(51 - Now Aaron)
Now Aaron = 51 - 2 now Aaron
3 Now Aaron = 51
Aaron = 17
Jack = 34
But Aaron is supposed to be twice Jacks age not half it.

Answer by Earlsdon(6294)

Try this.
A = Aaron's age now.
J = Jack's age now.
Write two equations from the problem description.
A = 2J Aaron is twice Jack's age.
(A-3) + (J-3) = 45 The sum of their ages 3 years ago = 45 Simplify.
A + J - 6 = 45 Add 6 to both sides.
A + J = 51 Substitute A = 2J
2J + J = 51
3J = 51 Divide both sides by 3.
J = 17 years.
A = 2J = 2(17) = 34 years.